OpenStudy (mackenzie2013):

Hans is building a fence to put around his rectangular garden. How many yards of fencing will Hans need to enclose the garden on all four sides? 10.5 yd 21 yd 7.5 yd 15 yd

OpenStudy (mackenzie2013):



OpenStudy (mackenzie2013):

@lacrosseplayer22

OpenStudy (anonymous):

so ur trying to find the area right?

OpenStudy (armyrangers):

No she should be looking for the perimeter

OpenStudy (anonymous):

No he is trying to find the perimeter. Because in a rectangle opposite sides are equal then you just do 6(2)+4.5(2)= 12+9=21 so thats your answer

OpenStudy (anonymous):

perimeter =2l+2w=12+9

OpenStudy (anonymous):

B 21 is the answer

OpenStudy (mackenzie2013):

Says B is incorrect.

OpenStudy (mackenzie2013):

Wouldn't you set up a proportion? Because you need to change the feet to yards?

OpenStudy (mackenzie2013):

Correct me if I'm wrong.

OpenStudy (armyrangers):

I missed that you are right

OpenStudy (anonymous):

4.5 ft=1.5 yard

OpenStudy (anonymous):

Oh wow we are all stupid.. yes you are right you shoud've changed feet to yards I missed that.

OpenStudy (anonymous):

3+12=15

OpenStudy (anonymous):

D sorry if u got the question wrong:(

OpenStudy (mackenzie2013):

Nah, its cool!